What is the total charge of all the electrons in 9.11 L of liquid water?
Answer #1

given

volume of the water v = 9.11 L

1 L of water contains 1000/18 = 55.6 moles

9.11 L water contains 9.11*55.6 = 506.5 moles

one mole contains 6.022*10^23 objects

so given water contains 506.5*6.022*10^23 = 3.05*10^26 water molecules

water has 10 electrons

number of electrons in 9.11 L of water is = 3.05*10^27

from the relation q = n*e

q = 3.05*10^27*(-1.6*10^-19 )= -4.88*10^8 C